OpenStudy (thickgirl_angel):

Given f(x)=4x+6 and g(x)=9x+9 then what is f(g(−6))? Given f(x)=3x+8 and g(x)=6x+6 then what is (f+g)(−1))? Given f(x)=3x+8 and g(x)=6x+6 then what is (f−g)(−1))?

OpenStudy (sshayer):

\[f \left\{ g \left( x \right) \right\}=4 g \left( x \right)+6=4\left( 9x+9 \right)+6=36x+36+6=36x+42\] now plug x=-6

OpenStudy (kittiwitti1):

\[f(g)\rightarrow f[g(x)]\]\[(f+g)(x)=f(x)+g(x)\]\[(f-g)(x)=f(x)-g(x)\]

OpenStudy (unklerhaukus):

Alternatively, find \(g(−6)\) first \[g(−6)=9(−6)+9=−45\], and plug it into \(f\) \[f(g(−6))=f(−45)=4(−45)+6\]

OpenStudy (unklerhaukus):

for the second part of the question use: \[(f+g)(x)=f(x)+g(x)\] and for the third part use: \[(f-g)(x)=f(x)-g(x)\]